How many kilograms of chocolates costing Php80.00 per kilogram must be mixed with 6 kilograms of chocolates that cost Php100.00 per kilogram to produce a mixture that can be sold for Php85.00 per kilogram?
Answer by Edwin McCravy(20055)

How many kilograms of chocolates costing Php80.00 per kilogram must be mixed with 6 kilograms of chocolates that cost Php100.00 per kilogram to produce a mixture that can be sold for Php85.00 per kilogram?

Let x = number of kilograms of lower priced chocolates to be used.

There are x kilograms of lower-priced chocolate

There are 6 kilograms of higher-priced chocolates

There are x+6 kilograms of mixed chocolates


Value of the lower-priced chocolates = (Php80.00)(x)
                  
Value of the higher priced chocolates  = (Php100.00)(6)

Value of the mixed chocolates = (Php85.00)(x+6)


Value                 Value                Value
  of                    of                   of
lower         +      higher        =       mixed  
priced               priced              chocolates 
chocolates          chocolates
                  
   80.00x     +      (100.00)(6)   =     (85.00)(x+6)

                        
               80x + 600 =  85(x + 6)
               80x + 600 =  85x + 510
              -85x       = -85x
              ------------------------
               -5x + 600 =        510
                   - 600         -600
               -----------------------
               -5x       =        -90
                       x = 18

18 kilograms of the lower-priced chocolates.

Edwin